OpenStudy (anonymous):

Find the nonpermissible replacement for the variable in this expression. 5x/ 6x+3 thats 5x over 6x+3

OpenStudy (anonymous):

a. -2 b. - 1/2 c. 0 d. 1/2

OpenStudy (solomonzelman):

I don't really get what the problem is like.

OpenStudy (solomonzelman):

just this, \(\large\color{black}{ \frac{\LARGE 5x}{\LARGE 6x+3} }\) ?

OpenStudy (solomonzelman):

okay, the non-permissible replacement is an x-value at which the expression is undefined.

OpenStudy (solomonzelman):

if the denominator is zero, then the expression is undefined. So let me ask you: When x equals what , does the denominator equal 0?

OpenStudy (solomonzelman):

(set the denominator equal to zero, and solve for x)

OpenStudy (anonymous):

i got 0

OpenStudy (solomonzelman):

it might seem as thought that you are trying to guess the answer. Please don't. You can ask if you don't get something, and I will attempt my best to explain this, but again please don't try to guess the.... So this is what we are doing here: \(\large\color{black}{ \frac{\LARGE 5x}{\LARGE 6x+3} }\). We know that; \(\large\color{black}{ \frac{\rm anything}{\LARGE 0}=\rm undefined }\) So, to find the values where \(\large\color{black}{ \frac{\LARGE 5x}{\LARGE 6x+3} }\) you need to equate \(\large\color{black}{ 6x+3 }\) to \(\large\color{black}{ 0 }\) . \(\large\color{black}{ 6x+3=0 }\) . can you solve for x?

OpenStudy (anonymous):

-1/2

OpenStudy (solomonzelman):

excuse me, to find the values *where \(\large\color{black}{ \frac{\LARGE 5x}{\LARGE 6x+3} }\) is undefined, you need to equate...

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

-1/2 is correct